OpenStudy (anonymous):

TRIG SUB Q:

OpenStudy (anonymous):

look it up in khan academy if they have itt

OpenStudy (anonymous):

Here is the problem: \[\int\limits_{}^{}\sqrt{16-x ^{2}}\]

OpenStudy (anonymous):

And here is ALL my work: http://imgur.com/K6tn1

OpenStudy (anonymous):

I think I pretty much have it solved, I just am unsure about that last conversion

OpenStudy (amistre64):

the minus suggests .... something sin^2 + cos^2 = 1 sin^2 = 1-cos^2 ; cos^2 = 1-sin^2 -------------------------------- tan^2 + 1 = sec; sec^2 = tan^2 - 1 some such conflagerations

OpenStudy (anonymous):

\[x=4 \sin (\theta )\]\[{dx}=4 \text{d$\theta $} \cos (\theta )\] \[\sqrt{16-(4 \sin (\theta ))^2}\] \[\sqrt{16-16 \sin ^2(\theta )}\] \[4 \sqrt{1-\sin ^2(\theta )}\] \[4 \sqrt{\cos ^2(\theta )}\] \[\int 4 \text{Cos}[\theta ] 4\text{Cos}[\theta ]\text{d$\theta $}\] \[16\int \text{Cos}[\theta ]^2\text{d$\theta $}\]